# PLeaS — Merging Models with Permutations and Least Squares

### v1

@@ Abstract
The democratization of machine learning systems has made the process of fine-tuning accessible to a large number of practitioners, leading to a wide range of open-source models fine-tuned on specialized tasks and datasets. Recent work has proposed to merge such models to combine their functionalities. However, prior approaches are restricted to models that are fine-tuned from the same base model. Furthermore, the final merged model is typically restricted to be of the same size as the original models. In this work, we propose a new two-step algorithm to merge models—termed PLeaS—which relaxes these constraints. First, leveraging the Permutation symmetries inherent in the two models, PLeaS partially matches nodes in each layer by maximizing alignment. Next, PLeaS computes the weights of the merged model as a layer-wise Least Squares solution to minimize the approximation error between the features of the merged model and the permuted features of the original models. into a single model of a desired size, even when the two original models are fine-tuned from different base models. We also present a variant of our method which can merge models without using data from the fine-tuning domains. We demonstrate our method to merge ResNet models trained with shared and different label spaces , and show that we can perform better than the state-of-the-art merging methods by 8 to 15 percentage points for the same target compute while merging models trained on DomainNet and fine-grained classification tasks.
@@ 1 Introduction
With the widespread democratization of machine learning, there has been a rapid increase in the availability of open-source models trained by the community on specific tasks and datasets. Such specialized models exhibit unique strengths and weaknesses. For example, Code Llama [25] (fine-tuned from Llama-2) is specialized for coding tasks, while Vicuña 1.3 [4] (fine-tuned from Llama-1) is specialized for chat. They have the same architecture but are fine-tuned starting from different pre-trained models: Llama-1 and Llama-2. Such diversity in the combination of pre-training data and fine-tuning tasks will increase as decentralized marketplaces for models become increasingly more common, e.g., [24], providing practitioners with more choices.
This presents an opportunity to combine such specialized models in order to create a single general-purpose model that can handle multiple tasks. Traditional approaches for combining trained models such as ensembling [8] or domain-specific mixture-of-experts (e.g.[13]) take a step towards this goal. However, these methods need to store all the component models at inference time, leading to an increased memory footprint. Practitioners with limited memory capacity cannot use such costly approaches with fixed memory footprints, especially when combining large models, deploying to resource-constrained environments, or for applications demanding a memory-performance trade-off.
To this end, recent works [32, 12, 35, 34] have proposed new algorithms tackling this problem of model merging. However, their scope is limited to merging models fine-tuned from the same pretrained model. Further, some recent works [28] also need access to the training data used to fine-tune the component models, which limits their applicability in situations where such data is not available due to, for example, privacy or legal reasons [7]. In this paper, we address the problem of merging models (sharing the same architecture) trained on different datasets starting from different initializations. This is motivated by prior work (e.g., [28, 33, 11]), which we compare with in Section 5 for merging ResNet models. Since transformer models exhibit more complex symmetries, we leave the task of merging such models for future work. To address the above-mentioned limitations of prior work in this space, we present PLeaS—an algorithm which adaptively merges models for different inference compute budgets, and can work without requiring data that the component models were fine-tuned on.
@@
PLeaS (short for Permutations and Least Squares) is a two-stage algorithm which works with models having the same architecture. The first step consists of matching features across the models. We harness the idea of permutation invariance in neural networks to find an appropriate pairing of features. Inspired by the Git Re-Basin [1] algorithm, which is designed for merging two models that are trained on the same data, we introduce a matching algorithm that finds permutations between similar features across models, while separating dissimilar features in the final merged model. This is critical when merging models trained on widely different tasks, since it prevents interference between features while still merging overlapping features, improving upon prior work such as ZipIt[28], which merges all the neurons of some layers of the two models. This also gives PLeaS fine-grained control over the width of each layer of the merged model. It can hence flexibly trade-off memory/compute and performance according to the deployment requirements.
@@
It has been observed that permutation matching alone suffers from significant performance loss when merging vastly different models, e.g. those trained on disparate data [33]. We hypothesize that while permuted features are powerful when ensembled, simply averaging the permuted weights degrades the features of the merged model. This results in the observed decline in performance.
Hence, in the second step of PLeaS, we solve a layer-wise Least Squares problem, so that each layer of the merged model mimics the permuted ensemble of features from the corresponding layer of the original models. This leads to better representations and superior down-stream performance.
@@
Apart from the target compute budget, PLeaS is hyperparameter free, making it easy for practitioners to use. A schematic of PLeaS is depicted in Fig. 1. We empirically demonstrate that PLeaS can outperform prior work in the challenging setting of merging differently initialized models which have been trained on different datasets.
@@
We merge ResNet models fine-tuned on different datasets in Sections 5.2 and 5.3, and find that PLeaS improves upon the state-of-the-art by 8 to 15% with the same merged model size. Our empirical results are on subsets of DomainNet, and on fine-grained classification tasks. Further, PLeaS, with significantly lower FLOPs, can approach the performance of ensemble methods in some cases (Section 5.5).
@@
The proposed approach can be seamlessly extended to the scenario where data from the fine-tuning domains is unavailable. We call this variant PLeaS-free. This variant uses data from publicly available datasets (like ImageNet) to merge models. We demonstrate in Section 5.4 that PLeaS-free can perform competitively to PLeaS which uses the actual data from the training domains of the component models. This is highly encouraging, as it demonstrates the applicability of PLeaS-free in scenarios where data from the training domains is unavailable due to privacy or commercial reasons.
@@
  • We generalize Git Re-Basin [1] to support partial merging of corresponding layers of two models (Section 4.1) and propose a strategy that automatically selects the target width of each layer of the merged model under a fixed FLOP budget (Section A.2). This gives the practitioner the freedom to choose the size of the final merged model as per resources available at inference. Investigating this tradeoff is one of the goals in this work, e.g., Fig. 3.
  • Motivated by the success of ensemble methods, we propose to assign weights to the merged model by solving a least squares problem attempting to mimic ensemble methods at each layer (Section 4.2). Ablation study for this step is in Fig. 3.
  • On a test-bed of multiple datasets (with both shared and different label spaces), we showcase that PLeaS outperforms recent merging methods by 8-15 percentage points.(Section 5) at the same model size. Further, PLeaS approaches the ensemble accuracy while using 40% fewer parameters. When no training data is available, PLeaS-free remains competitive to PLeaS (Fig. 4).
@@ 2 Related works
There has been growing interest recently in merging models with minimal data and compute overhead. Here, we focus on methods which merge models with the same architecture.
Merging models fine-tuned from the same initialization.
@@
Several methods aim to merge models in the weight space. Ilharco et al. [12] simply add up task vectors, the weight differences of fine-tuned models from the pretrained model, and demonstrate a strong baseline for merging fine-tuned models. Other approaches edit the task vectors based on magnitude of the weights [32, 36] to resolve interference while merging. Some methods aim to find layer-wise [35, 2] or parameter-wise [19] coefficients for merging different task vectors. However, such methods work with task vectors, assuming that the base pretrained model is shared across the fine-tuned models, and hence cannot be easily extended to settings where models are fine-tuned from different starting points. A different line of work [14] proposes layer wise distillation, aiming to minimize the sum of the $\ell_{2}$ distances between the activations of the merged model and the original models. However, naively applying this to models which are vastly different leads to performance degradation, as we show in Section 5. Further, these methods do not provide a way to control the size of the merged model. Although not designed for this scenario of merging fine-tunes of a common pre-trained model, PLeaS  still allows us to achieve significant performance gains when the merged network is slightly larger than the original model (e.g. by 20%) as demonstrated in Table 2.
Merging from two different initializations.
@@
We consider a less restrictive setting, where the models being merged can have different initializations. This has been studied in the literature, and existing works propose weight or activation matching algorithms for this task. Git Re-Basin [1] proposes an algorithm to compute permutation matrices to match the weights of the hidden layers of two or more neural networks. Yamada et al. [33] investigate the usage of permutations to merge models trained on different datasets, however, their study is limited to wide ResNet models on MNIST and CIFAR datasets. These permutation symmetries have also been studied in [20, 33, 5, 27]. Another recent work – ZipIt! [28], tackles a similar problem of merging models finetuned on different datasets from different initializations. ZipIt! proposes an alternative formulation for this task, allowing for feature matching across and within models and puts forth a greedy algorithm to optimize for this. ZipIt! also allows for merging up to some layers of the component models. While this can provide a knob for controlling the size-performance trade-off of the merged model, the empirical performance of their proposed scheme can be improved upon, as we show in Section 5.3. On the other hand, our work describes a merging formulation which is more expressive and allows for partial merges even within layers to minimize feature interference. Finally, He et al. [11] also propose a method to merge networks layer-wise in a progressive manner, which involves light-weight retraining. However, their method requires domain labeled data at both training and inference time, while we require unlabeled data only and also propose a method using no data at all.
Other merging paradigms
@@
Other model merging approaches include mixture of experts [26, 29], selecting experts using test data [18], and sparse expert ensembles [10]. These come with larger compute or memory over-heads, both at inference and training time.
@@ 3 Preliminaries
Notations. For simplicity, we describe our method for two $L$-layered MLPs. However, it can be readily extended to convolutional and residual networks, as we demonstrate in our experiments. Let $\Theta^{A}=\{W^{A}_{1},W^{A}_{2},\cdots,W^{A}_{L}\},\Theta^{B}=\{W^{B}_{1},W^{B}_{2},\cdots,W^{B}_{L}\}$ be the parameters of two MLPs $A,B$ having the same architecture. We omit the layer-wise bias here for simplicity. Let $z_{i}^{A},z_{i}^{B}$ denote the input activations to the $i$th layer of each network respectively, and $d_{i}$ denote the dimension of $z_{i}^{A},z_{i}^{B}$. We also define $Z_{i}^{A},Z_{i}^{B}\in\mathbb{R}^{d_{i}\times n}$ to be the activations of a batch of $n$ inputs. Note that $z_{1}^{A}=z_{1}^{B}=x$, and $z_{L+1}^{A}=y^{A},z_{L+1}^{B}=y^{B}$. Finally, let $\{W_{i}^{M}:i\in\{1,2,\cdots,L\}\}$ be the weights of the merged model. We allow the merged model to have varying widths (which can be different from the widths of the base model), depending on the memory and compute resources available.
@@
Ainsworth et al. [1] propose a method for computing the permutation matrices by directly optimizing the average similarity between permuted weights of model $B$ and the original weights of model $A$. This weight matching greedily finds a solution to the following sum of bilinear assignment problems,
@@ 4.1 Extending Git Re-Basin to partial merging
Note that in Git Re-Basin, two models are averaged (after permuting one model) and hence the dimension of the merged model is the same as the base models. However, when the networks $A,B$ are trained on different datasets, not all features might be compatible across models, and might interfere with each other if merged, leading to degraded performance. Further, those incompatible features might need to be retained in the merged model in order to make accurate predictions on both tasks. Merging all nodes in every layer discounts this possibility, leading to performance degradation, as we show in Fig. 4(b). To this end, we aim to merge features which are similar across the two models, while keeping those which are very different as separate features in the merged model. We hence propose a framework for partially merging model features by leveraging permutations.
@@
Figure 2: Partial merging with permutations: We show the construction of the $7\times 6$ weight matrix $W_{i}^{m}$ from two weights of size $5\times 4$. The merged inputs are copied and unpermuted to approximate the original inputs. Then we apply both weight matrices separately. Finally, we pair up the merged outputs and average the pairs. Since all operations used are linear, we can fuse them to construct $W_{i}^{m}$ using a single linear layer.
@@
This is simple to implement: retain the indices with the smallest $k_{i}$ distances between the (permuted) activations. For weight matching, we can retain the indices with the largest similarity between the (permuted) weights for each layer. The size of $W_{i}^{M}$ is then increased to $(2d_{i}-k_{i})\times(2d_{i+1}-k_{i+1})$ in exchange for improved performance. This partial merging scheme is illustrated in Fig. 2. Investigating this trade-off between the size and the performance of the merged model is one focus of this paper.
Note that the ratio of ${k_{i}}/{d_{i}}$ can be chosen to be different for each layer. In Section A.2, we introduce a scheme to find such a configuration of these ratios for a given target compute/memory budget $B$; this optimizes a proxy of the downstream performance without using any validation data from the target domain, and is used in all our experiments. The permutation matrices are computed using the weight matching strategy from Git Re-Basin. In Section 5.4, we compare this with using the activation matching strategy, which we call PLeaS-Act .

### v2

@@ Abstract
The democratization of machine learning systems has made the process of fine-tuning accessible to practitioners, leading to a wide range of open-source models fine-tuned on specialized tasks and datasets. Recent work has proposed to merge such models to combine their functionalities. However, prior approaches are usually restricted to models that are fine-tuned from the same base model. Furthermore, the final merged model is typically required to be of the same size as the original models. In this work, we propose a new two-step algorithm to merge models—termed PLeaS—which relaxes these constraints. First, leveraging the Permutation symmetries inherent in the two models, PLeaS partially matches nodes in each layer by maximizing alignment. Next, PLeaS computes the weights of the merged model as a layer-wise Least Squares solution to minimize the approximation error between the features of the merged model and the permuted features of the original models. PLeaS allows a practitioner to merge two models sharing the same architecture into a single performant model of a desired size, even when the two original models are fine-tuned from different base models. We also demonstrate how our method can be extended to address a challenging scenario where no data is available from the fine-tuning domains. We demonstrate our method to merge ResNet and ViT models trained with shared and different label spaces, and show improvement over the state-of-the-art merging methods of up to 15 percentage points for the same target compute while merging models trained on DomainNet and fine-grained classification tasks [^1] Code open-sourced at https://github.com/SewoongLab/PLeaS-Merging.
@@ 1 Introduction
With the widespread democratization of machine learning, there has been a rapid increase in the availability of open-source models trained by the community on specific tasks and datasets. Such specialized models exhibit unique strengths and weaknesses. For example, Code Llama [25] (fine-tuned from Llama-2) is specialized for coding, while Vicuña 1.3 [3] (fine-tuned from Llama-1) is specialized for chat. They have the same architecture but are fine-tuned starting from different pre-trained models: Llama-1 and Llama-2. Such diversity in the combination of pre-training data and fine-tuning tasks will only increase as decentralized marketplaces for models become increasingly more common, e.g., [24], providing practitioners with more options.
This presents an opportunity to combine such specialized models in order to create a single general-purpose model that can handle multiple tasks. Traditional approaches for combining trained models, such as ensembling [7] or domain-specific mixture-of-experts (e.g. [13]), take a step towards this goal. However, these methods need to store all the component models at inference time, leading to an increased memory footprint. Practitioners with limited memory capacity cannot use such approaches with high and fixed memory costs, especially when combining large models, deploying to resource-constrained environments, or for applications demanding a memory-performance trade-off.
To this end, recent works [33, 12, 36, 35] have proposed new algorithms tackling this problem of model merging. However, their scope is limited to merging models fine-tuned from the same pretrained model. Further, some recent works [28] also need access to the training data used to fine-tune the component models, which limits their applicability in situations where such data is not available due to, for example, privacy or legal reasons [5]. In this paper, we address the problem of merging models (sharing the same architecture) trained on different datasets starting from different initializations. This is motivated by prior work (e.g., [28, 34, 10]), which we compare with in Section 5 for merging ResNet and ViT models. To address the above limitations of prior work in this space, we present PLeaS—an algorithm which adaptively merges models for different inference compute budgets, and can work without using the fine-tuning data of the component models.
@@
PLeaS (short for Permutations and Least Squares) is a two-stage algorithm which works with models having the same architecture. The first step consists of matching features across the models. We harness the idea of permutation invariance in neural networks to find an appropriate pairing of features. Inspired by the Git Re-Basin [1] algorithm, which is designed for merging two models that are trained on the same data, we introduce a matching algorithm that finds permutations between similar features across models, while keeping dissimilar features separate in the final merged model. This is critical when merging models trained on widely different tasks, since it prevents interference between features while still merging overlapping features. This also gives PLeaS fine-grained control over the width of each layer of the merged model, improving performance over prior work such as ZipIt! [28]. PLeaS  can hence flexibly trade-off inference memory/compute and performance according to the deployment requirements.
@@
It has been observed that permutation matching alone suffers from significant performance loss when merging vastly different models, e.g. those trained on disparate data [34]. We hypothesize that while permuted features are powerful when ensembled, simply averaging the permuted weights degrades the features of the merged model. This results in the observed decline in performance. Hence, in the second step of PLeaS, we solve a layer-wise Least Squares problem, so that each layer of the merged model mimics the permuted ensemble of features from the corresponding layer of the original models. This produces better representations and superior down-stream performance.
@@
Apart from the target compute budget, PLeaS is hyperparameter free, making it easy for practitioners to use. A schematic of PLeaS is depicted in Fig. 1.
@@
We empirically demonstrate that PLeaS can outperform prior work in the challenging setting of merging differently initialized models which have been trained on different datasets. We merge ResNet and ViT models fine-tuned on different datasets in Secs. 5.2, 5.3 and 5.5, and find that PLeaS improves upon the state-of-the-art up to 15% with the same merged model size. Our empirical results are on subsets of DomainNet, and on fine-grained classification tasks. PLeaS can also approach the performance of ensemble methods with significantly lower FLOPs (Sec. 5.5).
@@
The proposed approach can be seamlessly extended to the scenario where data from the fine-tuning domains is unavailable. We call this variant PLeaS-free. This variant uses data from publicly available datasets (like ImageNet) to merge models. We demonstrate in Sec. 5.4 that PLeaS-free is competitive with PLeaS, which uses the actual data from the training domains of the component models. This is highly encouraging, as it demonstrates the applicability of PLeaS-free in scenarios where data from the training domains is unavailable due to privacy or commercial reasons.
@@
  • We generalize Git Re-Basin [1] to support partial merging of corresponding layers of two models (Sec. 4.1). This gives practitioners the freedom to choose the size of the final merged model as per resources available at inference. Investigating this tradeoff is one of the goals in this work, e.g., Fig. 3.
  • Motivated by the success of ensemble methods, we propose to assign weights to the merged model’s parameters by solving a least squares problem attempting to mimic ensemble methods at each layer (Sec. 4.2). Ablation study for this step is in Fig. 3.
  • On a test-bed of multiple datasets, we showcase that PLeaS outperforms recent merging methods up to 15 percentage points (Sec. 5) at the same model size. Further, PLeaS approaches the ensemble accuracy while using 40% fewer parameters. Finally, even with no data from the training domains, PLeaS-free remains competitive with PLeaS (Fig. 4).
@@ 2 Related works
There has been growing interest in merging models with minimal data and compute overhead. Here, we focus on methods which merge models with the same architecture.
@@
Merging models fine-tuned from the same initialization. Several methods aim to merge models in the weight space. Ilharco et al. [12] simply add up task vectors, the weight differences of fine-tuned models from the pretrained model, and demonstrate a strong baseline for merging fine-tuned models. Other approaches edit the task vectors based on magnitude of the weights [33, 37] to resolve interference while merging. Some methods aim to find layer-wise [36, 2] or parameter-wise [19] coefficients for merging different task vectors. However, methods that work with task vectors assume that the base pretrained model is shared across the fine-tuned models, and hence they cannot be easily extended to settings where models are fine-tuned from different starting points. A different line of work [14] proposes layer wise distillation, aiming to minimize the sum of the $\ell_{2}$ distances between the activations of the merged model and the original models. However, naively applying this to vastly different models leads to degraded performance, as we show in Sec. 5. Further, these methods do not provide a way to control the size of the merged model. Although not designed for this scenario of merging fine-tunes of a common pre-trained model, PLeaS still allows us to achieve significant performance gains when the merged network is slightly larger than the original model (e.g. by 20%) as demonstrated in Tab. 2.
@@
Merging from two different initializations. We consider a less restrictive setting, where the models being merged can have different initializations. This has been studied in the literature, and existing works propose weight or activation matching algorithms for this task. Git Re-Basin [1] proposes an algorithm to compute permutation matrices to match the weights of the hidden layers of two or more neural networks. Yamada et al. [34] investigate the usage of permutations to merge models trained on different datasets, however, their study is limited to wide ResNet models on MNIST and CIFAR datasets. These permutation symmetries have also been studied in [20, 34, 4, 27]. A recent work – MuDSC [32] leverages permutation symmetries both in weight space and activation space to merge models better, however, we show that PLeaS  outperforms this work empirically. Another recent work – ZipIt! [28], tackles a similar problem of merging models fine-tuned on different datasets from different initializations. This work also supports merging models partially by “zipping” some layers of the component models. While this can provide a knob for controlling the size-performance trade-off of the merged model, the empirical performance of their proposed scheme can be improved upon, as we show in Sec. 5.3. On the other hand, our work describes a merging formulation which is more expressive and allows for partial merges with expanded layers to minimize feature interference. Finally, [10] also proposes a method to merge networks layer-wise in a progressive manner, which involves light-weight retraining. However, their method requires domain labeled data at both training and inference time, while we only require unlabeled data and also propose a method using no data from the train domain at all.
@@
Other merging paradigms. Other model merging approaches include mixture of experts [26, 29], selecting experts using test data [18], and sparse expert ensembles [9]. These come with larger compute or memory over-heads, both at inference and training time.
@@ 3 Preliminaries
Notation. For simplicity, we describe our method for two $L$-layered MLPs. However, it can be readily extended to convolutional and residual networks, as we demonstrate in our experiments. Let $\Theta^{A}=\{W^{A}_{1},W^{A}_{2},\cdots,W^{A}_{L}\},\Theta^{B}=\{W^{B}_{1},W^{B}_{2},\cdots,W^{B}_{L}\}$ be the parameters of two MLPs $A,B$ having the same architecture. We omit the layer-wise bias here for simplicity. Let $z_{i}^{A},z_{i}^{B}$ denote the input activations to the $i$th layer of each network respectively, and $d_{i}$ denote the dimension of $z_{i}^{A},z_{i}^{B}$. We also define $Z_{i}^{A},Z_{i}^{B}\in\mathbb{R}^{d_{i}\times n}$ to be the activations of a batch of $n$ inputs. Note that $z_{1}^{A}=z_{1}^{B}=x$, and $z_{L+1}^{A}=y^{A},z_{L+1}^{B}=y^{B}$. Finally, let $\{W_{i}^{M}:i\in\{1,2,\cdots,L\}\}$ be the weights of the merged model. We allow the merged model to have varying widths (which can be different from the widths of the base model), depending on the inference resources available.
@@
Ainsworth et al. [1] propose to estimate the permutation matrices by directly optimizing the average similarity between the permuted weights of model $B$ and the original weights of model $A$. This weight matching greedily finds a solution to the following sum of bilinear assignment problems,
@@ 4.1 Extending Git Re-Basin to partial merging
Note that in Git Re-Basin, two models are averaged (after permuting one model) and hence the dimension of the parameters of the merged model is the same as the corresponding parameters of the base models. However, when the networks $A,B$ are trained on different datasets, not all features might be compatible across models. These features may interfere destructively if merged, leading to degraded performance. Further, these incompatible features may need to be retained in the merged model in order to make accurate predictions on both tasks. Merging all nodes in every layer discounts this possibility, leading to performance degradation, as we show in Fig. 4(b). To this end, we aim to merge features which are similar across the two models, while keeping those which are very different as separate features in the merged model. We hence propose a framework for partially merging model features by leveraging permutations.
@@
Figure 2: Partial merging with permutations: We show the construction of the $7\times 6$ weight matrix $W_{i}^{m}$ from two weights of size $5\times 4$ in the first step of PLeaS. The merged inputs are copied and unpermuted to approximate the original inputs. Then we apply both weight matrices separately. Finally, we pair up the merged outputs and average the pairs. Since all operations used are linear, we can fuse them to construct $W_{i}^{m}$ using a single linear layer.
@@
This is simple to implement: retain the indices with the smallest $k_{i}$ distances between the (permuted) activations. For weight matching, we can retain the indices with the largest similarity between the (permuted) weights for each layer. The size of $W_{i}^{M}$ is then increased to $(2d_{i}-k_{i})\times(2d_{i+1}-k_{i+1})$ in exchange for improved performance. This partial merging scheme is illustrated in Fig. 2. One goal of this paper is to investigate this trade-off between size and the performance of the merged model.
Note that the ratio ${k_{i}}/{d_{i}}$ can be chosen independently for each layer. In Sec. A.3, we propose a scheme to find a configuration of these ratios subject to a target compute/memory budget $B$; this optimizes a proxy of the downstream performance without using any validation data from the target domain, and is used in all our experiments. The permutation matrices $P_{i}$ are computed using the activation matching strategy from Git Re-Basin. In Sec. 5.4 and the Appendix, we compare this with using the weight matching strategy, which we call PLeaS-Weight .
We would like to emphasize that our partial merging formulation is different from ZipIt! [28], since we can assign any amount of compute between 1×1$\times$1 × and 2×2$\times$2 × (relative to the original layer’s compute) to each layer independently. ZipIt! on the other hand, assigns exactly 1×1$\times$1 × compute to a prefix of layers and 2×2$\times$2 × to the rest. In Sec. 5.3, we show that this flexibility in our formulation leads to better empirical performance.